OpenStudy (anonymous):

I'm not sure what type of science this is, so excuse me if this is the wrong section. What density is the kind of foam used in kids' trampoline? I need the info for a science fair project and I don't have a scale. I only need approximations. Also is it denser or less dense than printer paper? To clarify, the type of foam that shows up when you google "kids trampoline foam". Usually blue colored. Please help!! EDIT: If you can give me the technical term for the specific kind of foam, I think I can get the density from there. I don't know what it's called

OpenStudy (theeric):

This seems like a good section, but I really don't know! Good luck! As for which is more dense, consider holding two same-sized boxes: one box of foam and the other of paper. Which do you think would have more mass? I think paper.

OpenStudy (anonymous):

Even if I had a scale, the foam is thicker than the paper, so it'd be hard to calculate

OpenStudy (theeric):

Well, that's an interesting problem. You can't just compare their weights, because they have different sizes. However! You want to know about density, which is mass per volume, right? Well! You now need to find the mass and volume of each material. Think about paper, first. Weighing a sheet of paper might read "0" on the scale if the scale isn't sensitive enough. But, consider a \(stack\) of paper. Altogether, the stack will have the same density, because it's the same material. You might have 50 times the mass, but you also have 50 times the volume, so the density stays the same! Mathematically, \(\dfrac{50m}{50V}=\dfrac mV\). So you find the mass of all sheets on a scale, and then the volume. You can treat the stack like a box: height times width times thickness, or \(V=h\times w\times t\)